. A particle moves a fixed point O. The velocity v of the particle at time t s is given by v = (3t − 2)(t − 4) and s . A particle moves a fixed point O. The velocity v of the particle at time t s is given by

v = (3t − 2)(t − 4) and s = 8m when t = 1s, find:

(a) the initial velocity of the particle [1]

(b) the acceleration of the particle when t = 3s. [2]

(c) the values of t when the particle is at rest [2]

(d) the distance the particle is from O when t = 2s. [5]

(e) the distance travelled when t = 2s= 8m when t = 1s, find: (a) the initial velocity of the particle [1] (b) the acceleration of the particle when t = 3s. [2] (c) the values of t when the particle is at rest [2] (d) the distance the particle is from O when t = 2s. [5] (e) the distance travelled when t = 2s


1
Expert's answer
2021-10-31T18:15:57-0400

"v=(3t-2)(t-4)=3t^2-14t+8,"

"s(t)=\int vdt=t^3-7t^2+8t+C,"

"s(1)=1-7+8+C=2+C=8,\implies C=6,"

a)

"s(0)=6~m,"

b)

"a(t)=v'(t)=6t-14,"

"a(3)=18-14=4~\frac{m}{s^2},"

c)

"v=0: t=\frac 23~s~\text{or}~t=4~s,"

d)

"s(2)=8-28+16+6=2~m,"

e)

"s(1)=1-7+8+6=8~m."